# RIVCoin: an alternative, integrated, CeFi/DeFi-Vaulted Cryptocurrency

### v1

@@ 1.1 Towards a Decentralized Store of Value
  • Private Stablecoins, such as the JPM Coin.
This classification highlights two distinctive frames of reference in which stablecoins are labeled: the public vs. private nature of the stablecoin (i.e. whether the stablecoin can be publicly purchased and traded) and the intrinsic nature of the pegging mechanism (i.e. whether the stablecoin is backed by cash/cash equivalents, or if the peg is maintained through an algorithm).
Private Stablecoins are by nature centralized: there is a single entity that controls the emission of the currency. At the same time, Public Algorithmic Stablecoins heavily suffer from market downturns. This is the case of TerraUSD, collapsed in 2021 after a speculative attack.
We would like to highlight how the logic behind the general concept of stablecoin is similar to the one behind the mechanism of ETFs: instruments which exist to replicate the movements of a certain underlying. Unfortunately, if there is a disruptive detachment of correlation between the underlying to replicate and the basket used to achieve the performance of the replica, then the stablecoin will inevitably de-peg, in some cases causing the entire crash of the stablecoin.
Our solution is inscribed into the idea of providing a decentralized store of value, stabilized by a diversified portfolio of liquid assets: reserves are invested in funds holding a combination of traditional financial assets and cryptocurrencies, whose proceeds are distributed only to Liquidity Providers. At the same time, we are also focused on bringing an innovative cryptocurrency to Crypto Users, allowing them to pool the token to reap the rewards of the cryptocurrency markets without dealing with the complexity of the underlying protocols.
The cryptocurrency market still lacks a “gold standard-type” stabilized token traded in the market in the same way utility tokens are. It is fundamental to build tokens that are fully compliant with upcoming laws and regulations (ex. VARA recently official in Dubai, or Europe’s upcoming MICAR) and created and traded as tokens on the main blockchain platforms and ecosystems.
Still, such tokens have immense potential, since they can allow a whole potential market of Users to enter the cryptocurrency market: namely, Users that are now worried by the entry barrier that crypto requires.
2 Solution
In this section, we will discuss how we are planning to build the first cryptocurrency on the Cøsmos ecosystem stabilized by a diversified portfolio of CeFi and DeFi assets and that distributes to Liquidity Providers its overprice and/or part or all of its seignoirage created, using an innovative approach, which combines Centralized and Decentralized Finance.
Let us recall this simple, yet so profound principle: you cannot have a meaningful store of value backed by cryptocurrencies automatically shielded from the inherent volatility of the crypto space. This depends on the nature itself of cryptocurrencies: assets that only exists in the digital space, without a real physical counterpart.
Stablecoins have surely contributed to provide a fixed conversion between fiat and cryptocurrencies, via different pegging mechanisms. While fiat-backed stablecoins such as Tether and USDC have been used, still they rely on a central trusted party which backs every token with possibly enough sufficient cash equivalents. Crypto-Backed Stablecoins are inherently dangerous, since their peg relies on the stability of the cryptocurrency markets: while over-collateralization still mitigates the risks of default, the mechanism is still suboptimal from the beginning, since they protect the peg via assets falling in value when the market where the token is valued loses capitalization: it should be the other way around. Algorithmic stablecoins such as TerraUST, on the other end, tried to maintain the peg via the execution of an algorithm. The result has proven inadequate to the expectations of the holders, many of whom have lost all their money.
For this reason, we believe that, to guarantee a stable and fixed conversion between cryptocurrency and fiat currency, we must use a different approach. And the approach is to provide a token that is stabilized by a combination of both fiat and cryptocurrency.
While the combination of CeFi and DeFi assets will be the guaranty for the Users of the token, at the same time the decentralized nature of the token itself, with the possibility of being freely traded on many different blockchains, will remove the need to trust a third party: nevertheless, for the Authorities the existence of third parties also will be a further mechanism of control.
Actually, the solution we are building solves in a single cryptocurrency two different and related problems:
  1. 1.
    The existence of a decentralized token that is stabilized by real-world assets;
  2. 2.
    The existence of a cryptocurrency which allows non tech-savvy Users to enter the crypto space, while guaranteeing both liquidity and ease of use.
We are planning to solve this problem by providing a cryptocurrency called RIVCoin, which is minted by a Decentralized Autonomous Organization on the Cøsmos ecosystem if, and only if, a certain amount of real asset, as a reserve, is pledged (in a Vault deposit) as a minimum guarantee behind the issuing of the RIVCoin. The RIVCoin can be stored in a KYC/AML-compliant, digital non-custodial wallet called RIV Wallet. Afterwards, with the setup of an Osmosis Liquidity Pool, everyone will be free to purchase RIVCoin on any Cøsmos-compatible wallet.
The emission of RIVCoin will be subject to a very strict minting schedule: there will be a maximum coinbase amount of 99 billion RIVCoin (max cap), and each RIVCoin will be issued and made available to the market only after an equivalent investment has been acquired and vaulted as a pledge for all the users. In other words: the evaluated amount of reserves vaulted and pledged represents the quasi-full hedge of all the coins minted (both the primary minted and the ones in the liquidity pools).
A reasonable, fixed part of the total token supply will be pre-allocated to a separate Wallet and it will be used to compensate the team and the most relevant advisors adherent with the best incentives schemes. These tokens will be vested proportionally to the minting activity. The larger the minting the faster will be the premium that Users will release to this Wallet. This separate Wallet will be filled with a maximum cap of 4% of all the token supply. In other words, for every 100 EUR of collateral, the final token minted could be up to 104.
RIVCoin will be a fully decentralized cryptocurrency whose issuing will be carried out by a DAO (called RIV DAO). Users will be able to buy RIVCoin by downloading a digital, non-custodial Wallet called RIV Wallet, and performing the required Know Your Customer and Anti-Money Laundering procedures.
The fiat funds arising from the token sale will be securely stored in a Vault and will compose the underlying of the token. Reserves will be mostly used to purchase units of a Luxembourg SICAV RAIF (Reserved Alternative Investment Fund) in order to increase the reserves and/or the Liquidity Pool. A half of the invested reserves will go into a compartment invested in a model-driven, traditional quantitative strategy, not automated called A-RAIF, which diversifies into Fixed Income, thematic equity portfolios, and mutual/hedge funds, while the other half will be conveyed into another compartment called C-RAIF, which invests in hedge funds’ portfolios composed of outright positions on crypto assets, staking, and liquidity pools on the leading Decentralized Exchanges, through a portfolio of ISIN-coded funds.
Trusted Oracles will ensure the Proof of Assets, being able to prove, at every given moment, that all the assets that collateralize RIVCoin exist.
It is statistically unpredictable the case of all CeFi and DeFi reserves both down to zero, but we can argue that such a case can only happen either in case of extreme natural catastrophic events or in case of extreme fraudulent schemes (i.e. Madoff, FTX, etc.). While we are optimistic that the first case will not happen, for the second case we have set a blind mechanism of enlarged trust and checks, which drastically limits the possibility of fraud.
Cryptocurrency companies should never use their tokens as collateral. Instead, the focus should be on asset-referenced tokens and network-based connected liquidity pools that serve as “large reserves” to always ensure liquidity. The code is written to maintain an impeccably transparent and constant monitoring of the assets in the Vault, without permitting any distraction out of the investment framework agreed with the token’s owners.
We can say that RIVCoin is built on the three pillars of MPH: Meritocracy (Users gain the more they own coins and the more they make them work though the activities of pooling and staking), Pragmatism (Users final best interest is always covered by best effort coding and/or democratic decisions), and Honesty (Users have always clear rules to play with and no surprises can emerge which are not written in advance).
The way the code is written, paired with segregated and transparent accounts, does exclude any chance to use either fiat or cryptocurrencies to enter improper or fraudulent transactions. In other words, no bad surprise can happen. The price of the token can fluctuate due to demand and supply, but always in a transparent environment and within a strict legal frame.
The aim is to launch a cryptocurrency suitable to the highest number of Users. Traditional professionals are always eager to develop new skills, but they do not accept a superficial approach to risk management and to legal compliance. Hence, we developed a cryptocurrency which is suitable to CeFi and arouses the interest of DeFi Users. In fact, the code automatizes the pledge behind the token, net of fees quasi-full guaranteed, and at the same time, fixing the exposure at risk of the invested reserves, that consists of 50% traditional investments and 50% virtual asset investments (in the long run).
We will offer to our Users an innovative way of entering the crypto markets, which conjugates decentralization with a liquidity providing reward structure. This innovative way of thinking DeFi, namely a complement and not an adversary to Traditional Finance, will add value both for institutional players and smaller Users.
2.1 Business Strategy
We will now delve more in detail into our business strategy. As already mentioned before, RIVCoin is a cryptocurrency minted by a Decentralized Autonomous Organization, called RIV DAO. The issuing of RIVCoin follows a tight schedule: namely, a RIVCoin is minted for each new reference value unit of fiat currency pledged in the Vault. Since the invested reserves will increase the value of the Vault, the value of the RIVCoin should be, to some extent, affected by these increases.
When that happens, an opportunity to mint new currency emerges. In this case, a part of the increased liquidity will be cashed out and yield new RIVCoin, that will be democratically redistributed through all Liquidity Providers.
A change in the value of the reserves could (but must not) lead to a change in the same direction of the RIVCoin price, due to the perception of the Users. Reality is that the invested reserves fluctuate, and it is not due that by raising the level of the reserves the token necessarily appreciates its value. As a matter of fact, fluctuations of the RIVCoin cannot be prevented ex ante and, therefore, stabilization mechanisms need to be put in place as described in this Whitepaper.
If all these conditions hold, one (or more) RIVCoin are issued and sent to the User’s address. From that moment onwards, the sum will be sent into the Vault. The funds are distributed on two portfolios: the first portfolio will be allocated on traditional Fixed Income investments (such as high-yields bonds), Equity and Funds (mutual/hedge) through the A-RAIF, while the other 50% will fuel a portfolio that invests in staking and LP farming on the main decentralized protocols through the C-RAIF. All the investments are executed under a strict risk management control to assure the immediate liquidity at occurrence and convenience.
Part of the returns of the reserves will be distributed to the Liquidity Providers via incentive to the pool’s rewards, that will sum up to the rewards derived from trading fees.
We plan to launch RIVCoin starting from a Presale Phase, where selected Users will be able to acquire the token before the actual release. Following the Presale, the Primary Minting will be held: tokens will be sold at the reference price to all Users. We will then create a Liquidity Pool that will give holders the possibility of trading the token on the secondary market.
More in detail:
  • First Minting - Presale
    We will hold presales where a limited number of tokens will be released to selected Users at the reference price. To secure a spot in a presale, you will need to download RIV Wallet, perform all KYC/AML measures and digitally sign a Token Offering Request.
    The Token Offering Request is a formal document that states your intention to acquire a certain number of RIVCoin before a given date in the future. In the Token Offering Request, you must state the number of tokens you want to buy. After signing the document, you will still have 21 days if you wish to withdraw your order. The token offering will mint the number of tokens necessary to fill out the requests of all interested parties.
  • Development of RIV Wallet
    A digital non-custodial wallet has been developed: the RIV Wallet. This will allow Users to buy and sell RIVCoin. The wallet will be set up and managed by RIV Technologies FZE LTD, 100% subsidiary of RIV Capital Luxembourg, and will have all the functionalities of a known digital non-custodial Cøsmos wallet.
    RIV Wallet will be available on the main stores for Android and iOS as soon as possible. Everyone will be able to download the wallet from the website after the registration.
  • Token Minting - Primary Minting
    Primary Minting will then be open. Users will be able to buy RIVCoin through a predefined fiat on-ramp procedure. The minting price will initially be fixed to a reference value. The Primary Minting price will not necessarily remain the same, but it could fluctuate due to the different variables affecting the price (i.e., supply and demand, value of the reserves, etc.) subject to strict and rigorous protocols of risk monitoring, with the sole purpose of ensuring the stabilization of the RIVCoin.
  • Liquidity Pool Bootstrap
    We will set up a Liquidity Pool on Osmosis, that will allow Users to buy and sell RIVCoin in the secondary market against USDC and/or OSMO.
    The Pool could be incentivized with the usual rewards and/or the redistribution of the token’s overprice upon the activities of bonding. The choice of if and when to redistribute the token’s overprice is taken in the exclusive interest of the Users and the RIVCoin stability.
At regime, Users will be able to buy RIVCoin from both the Primary Market during the minting activity and, at a later stage, from the Secondary Market, buying it from the Liquidity Pool. When the max cap will be reached, Users will exchange the RIVCoin on a continuous secondary market.
While the Secondary Market will always be open, Primary Minting could be temporarily paused according to market conditions.
A sudden change in the RIVCoin price, or more in general a clear uptrend, signals an abrupt change in the token’s demand. The Primary Market price mitigates such trends, since it creates a natural arbitrage where Users could buy the token on the primary market and sell it for a profit in the Liquidity Pool, nullifying eventual token appreciations.
For example, if it is profitable to stimulate the demand for higher prices, a temporary halt to Primary Minting could be implemented. More in general, the Primary Minting price will be used for the sole purpose of stabilizing the price on the reference value.
Also, Primary Minting could be varied according to the reserves’ appreciation or depreciation. For example, if the reserves lose value, token mintings could be halted to avoid further dilution of value, or alternatively more RIVCoin could be minted at lower prices.
RIV Wallet will be available on the main App Stores for Android and iOS. We plan to distribute the token on the leading Decentralized Exchanges.
2.2 Technical Framing - The Cøsmos Ecosystem
We operated a careful choice of all the available blockchains and ecosystems, choosing the one that, in our idea, fits best with how we want to create and evolve the project. For this reason, we have chosen to build on the Cøsmos ecosystem.
We will now introduce Cøsmos, describing it and motivating the reasons why we chose to build on this innovative network of blockchains. If you already know Cøsmos, or you are simply not interested in these details, you can skip this section and go directly to section three.
Cøsmos [17] is a network of independent and interconnected blockchains. Together with Polkadot [23], it is one of the so-called “third generation” blockchains: to say, blockchains that aim to solve the scalability issues of Ethereum.
Ethereum is a unicum in the blockchain space: the first decentralized platform for the creation and execution of Smart Contracts, where no party has power over the others, has been one of the main innovations in the blockchain field, fueling a vast amount of new development. From Decentralized Finance to Non-Fungible Tokens, most of the blockchain applications have been built on Ethereum. It is the most used Smart Contract platform, with thousands of DApps built on it.
Still, Ethereum has a few fundamental issues: transaction fees are high, and it is poorly scalable. That has a direct effect on the applications that are built on it: the Ethereum blockchain can quickly become congested.
We believe that Cøsmos can solve all the issues of Ethereum, providing what is called the Internet of Blockchains. Building on Cøsmos will mean accessing an ecosystem of independent and trusted networks, easily exchanging value with one another.
The capability of Cøsmos blockchains to communicate with each other, using the IBC Protocol, together with the use of Tendermint [22] as a consensus algorithm and the low transaction fees, allows for a secure and resilient network.
Cøsmos is a vibrant ecosystem composed of different blockchains, called zones. Each Cøsmos zone has its peculiarities and aims to solve a different business problem. Use cases range from DeFi and Non-Fungible Tokens minting and issuance to more specialized services, such as building decentralized cloud networks.
Decentralized networks allow developers to build applications where Users can invest their liquidity on DeFi protocols based on other blockchains, such as Ethereum, using decentralized and secure bridges.
The Cøsmos Hub secures the Cøsmos network: a network of validators that use a Proof of Stake, Byzantine Fault Tolerant consensus algorithm, allowing them to stake their ATOM to gain rewards. It works like a hub, letting different blockchains communicate securely with one another.
Interoperability, scalability, and low transaction fees are the main reasons we have chosen to build on Cøsmos.
Users will have the possibility of accessing a wide range of DeFi instruments, not restricted to Cøsmos blockchains. Access to DeFi protocols on Ethereum, such as Yearn.Finance or Sushiswap pools will be possible.
RIVCoin is built upon a specific Cøsmos zone called Commercio.Network, which we will describe in the following paragraph.
2.3 Commercio.Network
The platform on which RIVCoin is built is a Cøsmos Zone called Commercio.Network [7]. It is a low-cost, scalable blockchain with Smart Contracts for natively minting Fungible and Non-Fungible Tokens. It enables many different use cases, such as the tokenization of physical or financial assets, the creation of Self-Sovereign Identities using W3C Standards (ex. DID, DID Document), and the digital signing of documents.
Rather than building on second-generation blockchains like Ethereum or Solana, we chose to build our solution on a blockchain developed explicitly for financial applications. The use of pre-built Smart Contracts makes it easier to safely mint and distribute tokens by implementing eIDAS standards.
Commercio.Network has, in fact, many interesting features, such as:
  • High Scalability
    Commercio.Network can manage thousands of transactions per second, providing a secure ecosystem for building and deploying blockchain applications. Scalability is one of the key issues when choosing a blockchain: second-generation blockchains like Ethereum would not have guaranteed us the minimum throughput necessary for our ecosystem to run.
  • Low Cost
    Transactions on Commercio.Network cost only 0.01€, plus VAT. Low transaction costs facilitate the exchange of value, allowing for a fairer ecosystem.
  • Sustainable
    Sustainability is one of their core values. Building on a sustainable blockchain is essential. Proof of Work-based blockchains have a higher environmental impact since miners need to solve the POW algorithm to be able to mine the subsequent block on the chain.
    Commercio.Network consumes in a year the same amount of computing power that a Proof of Work-based blockchain would consume in one second. That makes the network fully sustainable and able to function without having a high environmental impact.
  • Interoperable
    Being built on Cøsmos makes the blockchain fully interoperable. Commercio.Network can communicate with any other blockchain in the Cøsmos ecosystem using the Inter-Blockchain Communication protocol. Via Axelar or Gravity bridges, RIVCoin will be bridged in the future to the Ethereum Mainnet.
  • Legally Binding
    Commercio.Network follows eIDAS specifications: for this reason, transactions happening on this blockchain are legally binding. Everyone willing to buy RIVCoin must undergo a rigorous KYC/AML process. In this way, we are sure that you are who you say you are.
The use of pre-built Smart Contracts makes building applications on this network reasonably easy. Commercio.Network has, in fact, six Smart Contracts ready to use:
  • eMINT
    eMINT allows the minting of Fungible and Non-Fungible Tokens, making it possible to tokenize physical and non-physical assets.
  • eID
    eID allows the creation of digital, Self-Sovereign Identities compliant with all the directives of W3C, allowing to create and sign Verifiable Credentials and Verifiable Presentations.
  • eSignature
    eSignature allows you to sign PDF and XML documents digitally. Digital Signatures signed via this certificate are legally binding.
  • eDelivery
    eDelivery allows to notarize and store any document exchange between two parties. The document exchange has the same value as any paper document exchange.
  • eArchive
    eArchive allows you to store documents in a WORM (Write Once, Read Many) immutable object storage for decades at a minimum cost
  • eKYC
    eKYC allows you to easily handle KYC and AML procedures, simplifying all the necessary operations.
  • ePAY
    ePAY allows you to ask for a certified SEPA or credit card payment.
We are fully committed to developing blockchain ecosystems. For this reason, we chose to become a validator node of Commercio.Network, actively participating in the security of the protocol. Commercio.Network counts currently 96 different validators, among the most significant IT companies in Europe.
3 Governance
We plan to implement a mixed governance system to govern and direct the investments meaningfully. With this term, we are referring to a system where the regulated and authorized Asset Management Company determines the strategic asset allocation of the two macro portfolios that compose the Reserves investments: the CeFi and the DeFi one. In contrast, decisions on the percentage allocation between CeFi and DeFi can be affected (not necessarily modified) in a decentralized way.
The reserves investments will be implemented by applying pre-coded proprietary strategies, which have been fine-grained to be used in the cryptocurrency markets. RIV Capital experience in Asset Management and a consolidated realized track record have shaped the strategies to be effective in every market condition, showing a constant risk adjusted return in the top 1% of the world. Governance proposals will be blockchain-based and accessible to RIVCoin holders directly from their wallets.
Decisions affecting the rewards distribution via the incentivization of the Osmosis Liquidity Pool will be decided via internal voting. The amount of token distribution will be decided according to market conditions, performance of the reserves and need for stabilization for the price of RIVCoin. The amount of token distributed will be made available on public sources.
RIV Capital will periodically put suitable decisions to the vote: obviously, no technical decisions will be submitted with the potential consequence to compromise the stability of the asset allocation of the reserves. Only a subset of upgrade proposals will be subject to voting. RIVCoin holders can check the proposals subject to voting from their wallet’s “Governance” section. They will be able to analyze the proposals in detail and discuss them.
Voting will be open at a given date and time. Users who want to vote on the governance proposal will need to submit the vote within the specified timeframe (usually one week). After the temporal window has passed, Users will not be able to cast their votes.
4 Tokenomics
In this section, the Tokenomics of the project will be described. More in detail, the following topics will be discussed:
  • The minting schedule of RIVCoin
  • The strategic Asset Allocation between CeFi and DeFi portfolios of reserves
  • The potential rewards distribution to the Liquidity Providers
4.1 Definitions
We will start by providing the definitions of all the terms that will be used throughout the course of the section.
  • AML: Short Term for Anti-Money Laundering, refers to laws and regulations intended to stop criminals from disguising illegally obtained funds as legitimate income.
  • Centralized Exchange (CEX): An online platform, owned by one or more companies, that offers custody and exchange of Digital Assets.
  • Centralized Finance (CeFi): The ecosystem of companies and private institutions that offer financial services related to cryptocurrency and Digital Assets. Services can range from cryptocurrency custody and trading, clearing, loans and simplified access to Staking and Liquidity Pools.
  • Decentralized Autonomous Organization (DAO): An organization that operates on a decentralized blockchain network governed by smart contracts, that enables its members to make decisions through a consensus mechanism.
  • Decentralized Exchange (DEX): A protocol that allows decentralized exchange of Digital Assets through an automatic algorithm, called Automatic Market Maker. A Decentralized Exchange is composed by one or more Liquidity Pools where users can trade without relying on intermediaries such as brokerages, exchanges or banks.
  • Decentralized Finance (DeFi): The offering of financial services and products, such as lending, borrowing, trading, and investing, using decentralized platforms that run on blockchain technology.
  • Digital Asset: A term that defines any fungible or non-fungible token on a distributed ledger.
  • Farming: Locking tokens into high-yield strategies on DeFi protocols (ex. DEXes).
  • KYC: Short term for Know Your Customer, that is, all the procedures required by regulatory authorities that aim to identify a Users, prevent money laundering and other illegal activities.
  • Initial Liquidity Offering (ILO): The process of initial token distribution at the launch of a project, performed by adding Liquidity on a Decentralized Exchange.
  • ISIN: International Securities Identification Number, a 12-digit alphanumeric code that uniquely identifies a specific security.
  • Liquidity Pool: A Smart Contract that contains two or more digital assets, locked with the aim of providing liquidity to a Decentralized Exchange.
  • Minting: The process of emitting a cryptocurrency on a Distributed Ledger.
  • Net Asset Value (NAV): The underlying value of an investment fund’s assets less its liabilities, divided by the number of shares outstanding, expressed in fiat terms.
  • Presale: Any token offering that precedes the initial offering to the public.
  • Tokenomics: A term that indicates all the factors that define a token price, allocation and distribution to Users, including any factor that can change the token supply and/or demand (such for example token burn or incentive mechanisms).
4.2 RIVCoin
The ecosystem is fueled by RIVCoin: a cryptocurrency minted by RIV DAO that is designed to have a value which fluctuates around a stable reference. RIVCoin has a fixed cap of 99 billion tokens. The initial selling price of the token in the primary market will be chosen according to the main fiat currencies. If market conditions vary abruptly, such as if a sudden increase in demand doubles the price of the token, RIVCoin’s reference value could change. The token minting schedule will reflect such change.
Note that the reference value does not necessarily coincide with neither the primary nor the secondary minting price. To avoid price speculations different from normal supply and demand equilibriums, all main numbers will be publicly disclosed in order for the Users to be able to judge at any time.
  1. 1.
    Initial Liquidity Offering (ILO) Presale
    During the ILO Presale phase, we will collect funds from selected Users that have expressed their interest in the project. Users will acquire the token at the starting reference value.
    To secure a spot in the presale, Users need to complete all the required KYC/AML procedures. Orders will be filled using the first come, first served principle: whoever can reserve a spot first will be able to receive tokens before the others.
  2. 2.
    DEX Sale + Residual ILO
    Following the Presale Phase, a Liquidity Pool will be built. Users will be able to pool their funds for rewards directly from their Wallet to the Liquidity Pool, and freely trade their RIVCoin against USDC (as well as OSMO at the soonest). Users can decide to pool their tokens in the Osmosis Liquidity Pool and receive the reward distribution given by the transaction fees together with the pool incentives.
    After the Liquidity Pool launch, we still plan to hold a certain number of token issuing with primary token emissions (i.e. minting new tokens): both primary and secondary markets will be available in this phase.
  3. 3.
    DEX Sale Only
    After all the established RIVCoin will be issued and the maximum number of tokens is reached, the only way for Users to acquire RIVCoin will be to buy them on the secondary market, on the Liquidity Pool on Osmosis, or any other Decentralized Exchange that supports it.
We forecasted a long primary dealing minting to allow the Community to study and understand the token and familiarize with it. The period to release all the RIVCoin available could be foreseen as a decade. Within that period, it is reasonable to expect adjustments and fine tuning of the model. By that time in the long run, only a listed, secondary market will be available and all the RIVCoin will remain fully collateralized by real assets (Equity, Bonds, Funds).
We plan to open a limited number of presales and allot a minimum purchase for entering the presale. Presales will be announced on RIVCoin’s website and on Twitter, as well as on any social media which can be used in full compliance.
4.3 Reserves Allocation
The Vault’s reserves are granularly invested in a wide range of financial products, ranging from high-yield Fixed Income securities to DeFi staking and liquidity pools. On the basis of the model, we expect a close to zero delinquency rate. This makes the token fully stabilized and shields the Users from the high downward volatility usually experienced in cryptocurrency markets.
More in detail, the strategic asset allocation that backs RIVCoin is summarized as follows. In the long run:
  • 50% (or more) of the reserves will be invested in a model-driven, traditional quantitative strategy that diversifies into Fixed Income, thematic equity portfolios, and mutual/hedge funds.
  • Up to 50% of the reserves will be invested in portfolios composed of outright positions on crypto assets, staking, and liquidity pools on the leading Decentralized Exchanges.
Diversification of the vaulted assets ensure that, if a black swan event happens, Users are shielded from catastrophic losses of the token’s reserves. Of course, volatility can always increase, but the expectation is that the risk/reward will be on an optimal point, marking an acceptable position given the expected return associated to that variability in prices.
RIVCoin is born as a cryptocurrency fully stabilized by assets. As such, its value should never eventually slow below the stabilized Net Asset Value price of the invested reserves, since natural arbitrage mechanisms will push the price back to it.
No leverage will be ever implemented using the tokens of the Users as a collateral, no matter the haircut granted by the banking system. The code is written exclusively to build up the reserves, properly pledged in favour of all Users holding the tokens.
Portfolio allocation will be decided via selected strategies that will strike a balance between risk and reward, always in the exclusive interest of the Users and following the highest standard practice. The staking and LP farming portfolios of the reserves will be diversified enough to minimize market risk: Liquidity Pools will be chosen on the leading Decentralized Exchanges, from Osmosis to Uniswap.
4.4 Token Allocation
The allocation of tokens will follow the principles of equality and fairness. Tokens will only be issued after a user’s purchase.
Also, the issuing mechanism is built such that each User, by buying RIVCoin and contributing to fill the Vault with new reserves, enables the issuing of other RIVCoin, thus actually acting as miner of new currency. The miner and User at time $t_{-1}$ is RIV Capital or one of its companies within the Group investing the very first seed. The users at time $t_{0}$ are the miners of the RIVCoin minted in $t_{1}$, the Users in $t_{1}$ becomes the miners of the coins minted in $t_{2}$, and so on. Every User purchases coins already stabilized, vaulted and secured.
We believe that enabling a secondary market for RIVCoin will lead to a fairer network, allowing Users to gain full control of their investment. This contrasts with black boxes where Users invest their money without any possibility to use and unwind it.
The goal is to create a Gold Standard for the world of cryptocurrencies, as the token is immediately convertible into assets already valued, be they CeFi or DeFi, in full credibility of the coin itself being convertible in last resort, being the Vault the “Seller of Last Resort” (SoLR).
The following rules will be strictly applied without any exception:
  • The portfolio allocation is decided within the regulated and vigilated Asset Management Company managing the SICAV RAIF.
  • The cryptocurrency allocation of the reserves cannot be over 50% of the total reserves. This will prevent overexposure to cryptocurrency.
  • In case of a depreciation of the RIVCoin price of more the 30% of its value in a week’s timeframe, the Liquidity Pool contract on Osmosis could be paused (Safeguard Policy). Such pausing will have the sole purpose of preventing an excessive depreciation of the RIVCoin, safeguarding the peg to the reference value.
  • In case of overshooting of the RIVCoin, the governance could decide to intervene with open market measures buying/selling RIVCoin.
4.5 Rewards, NAV and Applied Fees
At the end of each month, if the vaulted reserves have increased in value and there is a perceived changed value of RIVCoin, new tokens could eventually be distributed to RIVCoin Liquidity Providers, always with the aim to stabilize the price around the reference value. In this case, RIVCoin holders will then receive rewards in token terms, that will be calculated from the amount of RIVCoin LP Tokens they have in their Wallet directly from the Osmosis protocol. Users and analysts will always be able to verify the performance of the reserves by analyzing the RAIF’s monthly NAV chart and the DAO’s monthly yield. Rewards made available to Liquidity Providers will vary depending on the performance of CeFi and DeFi portfolios.
Liquidity Rewards will be paid in RIVCoin, as the standard practice. Liquidity Providers will have the opportunity to cash out and/or to reinvest the rewards. Rewards will be paid net of performance fees.
A Swap Fee (transaction fee) will be applied when the token is purchased and sold, of the amount of 0.03%. The Entry Fee (paid when buying RIVCoin in the primary market) will be 5%.
[…]

### v2

@@ 1.1 Background
  • Private Stablecoins, such as JPM Coin.
This classification highlights two fundamental axes: the public versus private nature of issuance, and the mechanism of value stabilization–whether through hard collateralization or algorithmic control. Each model involves trade-offs among decentralization, resilience, and regulatory exposure.
Public reserve-backed stablecoins depend on centralized custodians for collateral management, introducing custodial and regulatory risk. Algorithmic stablecoins, in contrast, attempt to maintain pegs through supply manipulation, but remain highly vulnerable to market shocks, as seen in the collapse of TerraUSD.
1.2 Toward a Reserve-Backed Digital Economy
Although stablecoins have provided a degree of price stability within crypto markets, they inherit many of the structural vulnerabilities of the fiat systems they mirror. Both fiat currencies and stablecoins ultimately lack a key property that underpinned historical monetary systems like the gold standard: intrinsic reserve backing. Under the gold standard, monetary value was directly tied to verifiable, tangible assets. The physical limitation of gold supply enforced fiscal discipline, anchoring currency stability across generations.
Modern fiat currencies are primarily supported by debt issuance and dynamic monetary policy, rather than by tangible reserves. Central banks manage fiat stability through interest rates and open market operations, aiming to control inflation and promote economic growth. They are, in essence, algorithmic stablecoins underwritten by economic output and sovereign enforcement, rendering them more resilient, yet fundamentally prone to the same structural instabilities. However, this system inherently depends on the perpetual expansion of liabilities – governments continually issue new debt to finance spending, while the money supply grows to accommodate these obligations. Over time, this dynamic has led to cycles of inflation, monetary devaluation, and systemic financial fragility.
For a currency, digital or otherwise, to serve as a true store of value and unit of account, it must break free from debt-dependence and/or artificial pegs. It must be supported instead by a growing, diversified base of real reserves that reflect genuine productive value.
Stablecoins have made important contributions by bridging fiat and crypto markets, but they are transitional instruments rather than final solutions. We believe the next evolution requires returning to first principles: a monetary system where value is directly anchored in real-world assets, professionally managed and dynamically adapted to changing economic conditions, but without relying on perpetual debt expansion.
This is the philosophy behind RIV Coin: a currency backed by an actively managed treasury (“The Vault”), composed of tangible, productive assets, rather than debt instruments or unsupported issuance mechanisms. Through Proof of Integrity, regular audits, and a sustainable treasury strategy, RIV Coin aims to provide a reliable, scalable, and inflation-resilient monetary alternative for the digital age.
2 RIV Coin, a reserve backed cryptocurrency
Our purpose is to create a cryptocurrency, with emphasis on currency, a word which gets its origin “from Medieval Latin currentia, literally: a flowing, from Latin currere to run, flow”, and literally meaning “medium of exchange” [currencydefinition].
We aim to develop a stable medium of exchange backed by an actively managed reserve of real assets, combining transparent, professional treasury management with decentralized security models.
We will accomplish this through the introduction of RIV Coin, a native Layer 1 cryptocurrency backed by a diversified reserve portfolio–hereinafter referred to as the Vault–which holds both traditional financial (CeFi) assets and decentralized finance (DeFi) positions. In the system we are building, the Vault guarantees the total number of coins minted and, as a last resort, it can always be sold to reimburse part/all the coins in circulation. the Vault can fluctuate in its value, but it cannot be used for anything else except for guaranteeing the minimum intrinsic value of all the coins minted.
The Vault’s CeFi holdings are initially periodically, then in real time, audited and notarized by professional independent third parties (typically fund administrators and/or auditors), while the DeFi positions remain transparently verifiable on-chain. The combined portfolio undergoes continuous proof generation, ensuring that the system can attest to its Net Asset Value (NAV) without exposing sensitive operational data. This structure fosters verifiable trust, protects against information asymmetry, and mitigates exploitative practices such as front-running.
RIV Chain, the native network supporting RIV Coin, utilizes a Proof-of-Stake consensus mechanism in which validators and their delegators secure the chain and earn yield proportional to the Vault’s growth.
The seignorage, defined as the return generated by the reserves pledged within the Vault, will be used for a democratic distribution of wealth, subject to explicit actions to be implemented by the Users. In contrast to traditional economies where seigniorage often drives “bad inflation” (as it makes governments to be self confident and print money beyond real growth) through uncontrolled monetary expansion, RIV Coin’s model redistributes the return transparently and sustainably. Growth from the underlying assets flows back to validators and their delegators, aligning incentives while promoting “good inflation” – an increase in value directly linked to productive asset management rather than unchecked monetary supply.
2.1 Growing Reserves
The long-term stability of RIV Coin is underpinned by the Vault: a diversified reserve invested in actively managed traditional (CeFi) and decentralized finance (DeFi) assets. The Vault is designed not merely to protect the value of RIV Coin, but to grow it sustainably over time through long-term investment strategies, low turnover, high risk-adjusted returns.
Reserves are continuously expanded through two primary mechanisms:
  • Primary Reserve Growth: New capital inflows are generated through non-inflationary minting mechanisms, where newly issued RIV Coins are backed 1:1 by fresh fiat or crypto assets pledged into the Vault.
  • Organic Investment Returns: the Vault’s assets are strategically deployed across diversified financial products–in a long-term, granular asset allocation. This produces organic return that grows the reserve base without requiring new issuance.
To safeguard the integrity and sustainable expansion of reserves, the Vault employs several critical design principles:
  • Zero-Knowledge Transparency: the Vault’s Net Asset Value (NAV) is periodically validated through zero-knowledge proofs, maintaining performance transparency without revealing sensitive portfolio details. This structure protects against front-running and other exploitative behaviors.
  • Independent Auditing: Regular third-party audits independently verify the composition and value of the Vault, ensuring user confidence in the reserve’s true backing.
  • Resilience through Diversification: Assets are granularly diversified across sectors and geographies to minimize systemic risk. Even in the event of a black swan market disruption, the impact on the Vault’s overall value is highly mitigated, and the overshooting effects are neutralized in the short run.
[…]